### v3

Abstract
Large language models (LLMs), with their powerful generative capabilities and vast knowledge, empower various tasks in everyday life. However, these abilities are primarily concentrated in high-resource languages, leaving low-resource languages with weaker generative capabilities and relatively limited knowledge. Enhancing the multilingual capabilities of LLMs is therefore crucial for serving over 100 linguistic communities worldwide. An intuitive approach to enhance the multilingual capabilities would be to construct instruction data for various languages, but constructing instruction data for over 100 languages is prohibitively costly. In this paper, we introduce BayLing 2, which efficiently transfers generative capabilities and knowledge from high-resource languages to low-resource languages through language alignment. To achieve this, we constructed a dataset of 3.2 million instructions, comprising high-resource language instructions (Chinese and English) and cross-lingual instructions for 100+ languages and performed instruction tuning based on the dataset to facilitate the capability transfer between languages. Using Llama as the foundation model, we developed BayLing-2-7B, BayLing-2-13B, and BayLing-2-8B, and conducted a comprehensive evaluation of BayLing. For multilingual translation across 100+ languages, BayLing shows superior performance compared to open-source models of similar scale. For multilingual knowledge and understanding benchmarks, BayLing achieves significant improvements across over 20 low-resource languages, demonstrating its capability of effective knowledge transfer from high-resource to low-resource languages. Furthermore, results on English benchmarks indicate that BayLing maintains high performance in high-resource languages while enhancing the performance in low-resource languages. Demo [^1] Demo: http://nlp.ict.ac.cn/bayling/demo  3Homepage: http://nlp.ict.ac.cn/bayling, homepage22footnotemark: 2, code [^3] Code: https://github.com/ictnlp/BayLing   5Models: BayLing-2-7B, BayLing-2-13B, BayLing-2-8B and models44footnotemark: 4of BayLing are available.
1 Introduction
In recent years, the field of natural language processing (NLP) has witnessed a significant surge in the development and utilization of large language models (LLMs) (OpenAI, 2022, 2023). Equipped with rich knowledge, strong generative capabilities, and diverse instruction-following abilities, LLMs empower various specific tasks such as translation, summarization, chat and question answering, seamlessly integratd into everyday life.
However, a significant portion of these potent capabilities is primarily concentrated in English, stemming from the fact that high-resource languages, with English as the representative, occupy over 90% of the pre-training and fine-tuning corpora (Chang et al., 2023; Touvron et al., 2023; Xu et al., 2024). This results in issues such as lack of knowledge and lower generative capabilities in many other low-resource languages (Nguyen et al., 2023a; Alabi et al., 2022). It is imperative to recognize that linguistic diversity is a fundamental aspect of human communication, with over 7000 spoken languages worldwide, and more than 200 of them are writable (Chang et al., 2023). The accelerating force of globalization underscores the importance of leveraging LLMs to serve diverse linguistic communities.
Enhancing the multilingual capabilities of LLMs is not a trivial task. The intuitive approach is to construct instruction data for various languages to enhance the LLM’s ability to follow instructions and generate responses across different languages (Zeng et al., 2023). However, given the extremely limited instruction data available for some low-resource languages and the prohibitive manual efforts required to construct instructions for over 100 languages, this approach becomes impractical. Therefore, exploring more efficient approaches to improving the performance of LLMs across diverse languages remains an area for further investigation.
On these grounds, we developed BayLing 2, a multilingual LLM, which transfers knowledge, generative capability and instruction-following ability from high-resource to low-resource languages through fine-tuning LLMs on cross-lingual tasks. Previously, BayLing 1 successfully explored transferring English knowledge and capabilities to Chinese through cross-lingual alignment (Zhang et al., 2023). Building upon BayLing 1, BayLing 2 extends language alignment to multilingual settings, particularly between high-resource and low-resource languages, leading to a multilingual LLM. The fine-tuning corpus of BayLing 2 primarily consists of Chinese and English instructions, supplemented with rich cross-lingual task instructions between Chinese/English and over 100 other languages, facilitating the capability transfer across languages.
Based on foundational models Llama-2-7B-Chat, Llama-2-13B-Chat, and Llama-3-8B-Instruct, we developed BayLing-2-7B, BayLing-2-13B, and BayLing-2-8B through the proposed efficient language alignment. We conducted a comprehensive evaluation of BayLing’s performance on both multilingual and general tasks and assessed the quality of language alignment through multilingual translation using the Flores-101 and WMT22 benchmarks. BayLing showed superior translation performance across more than 100 languages, achieving the best results among open-source models of comparable scale. We further evaluated BayLing’s multilingual knowledge and generative capabilities using benchmarks including Belebele, Multilingual HellaSwag, XNLI, and Multilingual ARC. The results indicated significant performance improvements across more than 20 low-resource languages, such as Bambara, Luganda, Swahili, and Zulu. This demonstrates effective knowledge and generative capability transfer from high-resource to low-resource languages. Additionally, we evaluated BayLing on various general benchmarks (primarily in English), finding that the language alignment had minimal impacts on BayLing’s performance in high-resource languages.
By further analyzing the experimental results, we get the following findings:
  • By fine-tuning on high-resource language instructions and cross-lingual instructions, LLM can transfer knowledge and generative capabilities from high-resource languages to low-resource languages, thereby facilitating multilingual interaction.
  • Cross-lingual instructions, such as interactive translation and multilingual translation, can efficiently enhance the language alignment within LLM, thereby improving translation performance.
  • Fine-tuning LLM solely on high-resource language instructions will involve inter-language conflicts and significantly impair the multilingual capabilities of LLM, especially on the low-resource languages. Beside high-resource language instructions, introducing cross-lingual instructions can effectively solve this issue.
Figure 1: Overview of BayLing 2. BayLing 2 is a multilingual LLM with efficient language alignment. BayLing 2 designates Chinese and English, two high-resource languages, as pivot languages and applies cross-lingual tasks to align 100+ languages to these pivot languages, which facilitates the capabilities transfer from high-resource languages to low-resource languages. During inference, BayLing 2 is capable of high-quality interaction across multiple languages.
2 Related Work
Multilingual LLMs, with their capability to handle and produce content in multiple languages simultaneously, hold promise for serving diverse linguistic communities. Foundational models, such as Llama (Touvron et al., 2023), GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2022), OPT (Zhang et al., 2022) and GLM (Du et al., 2022), are pretrained on corpora sourced from the web and books, which often encompass multiple languages. However, the distribution of languages in these corpora is notably imbalanced. Specifically, a few high-resource languages dominate a significant portion of the corpus, while a vast number of low-resource languages occupy only a small fraction (Touvron et al., 2023). This leads to performance variations across different languages (Ojo and Ogueji, 2023; Nguyen et al., 2023a). Moreover, subsequent supervised fine-tuning on English-centric instruction data exacerbates the issue of language imbalance (Lai et al., 2023), rendering LLMs lower interactive capability with low-resource languages.
Current approaches mainly fall into two categories: continual pretraining and supervised fine-tuning. With continual pretraining, some works focus on continuously pretraining foundational models using multilingual corpora to enhance their multilingual capabilities (Nguyen et al., 2023b; Lai et al., 2023; Ke et al., 2023; Gupta et al., 2023). These approaches effectively supplements LLMs with multilingual knowledge and generation abilities. However, continual pretraining often relies on large amounts of multilingual data, and thereby the costs associated with data collection and training are significant (Nguyen et al., 2023b; Liu et al., 2024). Moreover, there is a risk of catastrophic forgetting with continual pretraining, which may compromise the performance of the foundational model on high-resource languages (Li et al., 2024). Additionally, since the pretraining corpora of foundational models are often close-sourced, it is challenging to maintain the same distribution between the continual pretraining data and the pretraining data, which may lead to conflicting knowledge and potential hallucinations.
For supervised fine-tuning, existing methods attempt to manually annotate multilingual instructions to activate LLMs’ ability for multilingual interaction (Eisenschlos et al., 2020; Alabi et al., 2022; Lai et al., 2023; Wang et al., 2024; Shaham et al., 2024). This approach often relies on manually annotation and overlooks leveraging the capabilities of foundational models in high-resource languages as well as the generalization ability of LLMs. To address this, BayLing 2 attempts to enhance the multilingual capabilities of LLMs in a more efficient manner. The instruction dataset of BayLing 2 comprises instructions in both high-resource languages and cross-lingual instructions. The instructions in high-resource languages are designed to activate LLMs’ instruction-following capability, while cross-lingual instructions aim to facilitate multilingual alignment of LLMs, thereby transferring knowledge, instruction-following, and generation abilities from high-resource languages to low-resource languages.
3 BayLing 2
We introduce BayLing 2, a LLM equipped with enhanced multilingual capabilities through efficient language alignment. Building upon open-source foundational models, BayLing 2 endeavors to explore an efficient and cost-effective approach to enhance the multilingual capabilities, thereby addressing the demands for multilingual interaction.
3.1 Multilingual Alignments with Cross-lingual Tasks
During the pre-training stage, the distribution of languages in the corpus is highly imbalanced. For instance, English, being a high-resource language, accounts for over 90% of the corpus, while low-resource languages such as Sinhalese, Marathi, and Macedonian collectively comprise less than 1% of the corpus. Naturally, foundational models trained on such language-imbalanced corpus exhibit superior performance on English compared to low-resource languages. Previous studies have often noted that due to the generalization capability, LLMs also demonstrate a certain advantage on those languages within the same language family as English. Naturally, aligning low-resource languages with high-resource languages already mastered by LLMs allows us to transfer the knowledge and generation capabilities of LLMs from high-resource languages to other languages efficiently, thereby enhancing the multilingual capabilities of LLMs.
We employ cross-lingual tasks to align low-resource languages with high-resource languages, thereby achieving multilingual alignment. Fortunately, translation tasks naturally serve as well-defined cross-lingual tasks, demanding outputs that maintain consistent meanings with inputs while differing in language. More importantly, translation tasks boast abundant high-quality parallel corpus across diverse domains, thus laying the groundwork for the efficient achievement of language alignment.
Figure 2: Language distribution of instruction dataset.
Figure 3: Distribution of instruction categories, including Chinese, English and cross-lingual instructions.
Figure 4: Distribution of the tokens number involved in each instruction.
Specifically, we designate Chinese and English, two high-resource languages, as the pivot language, and align over 100 other languages to Chinese and English using translation instructions. Following this idea, we construct the instruction dataset for BayLing 2, comprising Chinese instructions, English instructions and cross-lingual instructions, as shown in Figure 1. The instruction dataset contains a total of 3.2 million instructions (1471 million tokens), with the distribution of Chinese, English and cross-lingual instructions illustrated in Figure 4. Notably, the cross-lingual instructions in BayLing 2 involve interactive translation, constrained translation, document-level translation and single-sentence translation tasks across over 100 languages. The language distribution of the instruction dataset is shown in Figure 2. The distribution of topics covered in the proposed instruction dataset is illustrated in Figure 4, where instructions primarily sourced from news corpora ensure data quality and security. Overall, BayLing 2 is fine-tuned on 3.2 million instructions covering 100+ languages, achieving multilingual alignment and thereby transferring knowledge and generation capabilities from high-resource languages to low-resource languages.
3.2 Training
Figure 5: Training loss curve of BayLing-2-8B.
Using Llama-2-7B-Chat, Llama-2-13B-Chat and Llama-3-8B-Instruct as foundational models, We fine-tune BayLing-2-7B, BayLing-2-13B and BayLing-2-8B respectively on the instruction dataset proposed in Section 3.1. We fine-tune BayLing 2 on 8 NVIDIA A800 80G GPUs for 3 epochs, using a global batch size of 128, learning rate of 2e-5 and weight decay of 0.0. Note that we apply learning rate of 2e-6 for BayLing-2-8B. We employ DeepSpeed (Rasley et al., 2020) and Gradient Checkpointing (Chen et al., 2016) techniques to optimize memory consumption. The training loss curve of BayLing-2-8B is depicted in Figure 5.
4 Evaluation
In this section, we comprehensively evaluate the performance of BayLing-2-7B, BayLing-2-13B and BayLing-2-8B on multilingual tasks and general tasks respectively.
4.1 Multilingual Capability
BayLing’s multilingual capabilities are primarily manifested in two aspects: multilingual translation and multilingual interaction. Multilingual translation aims to accomplish translation between different languages, which can be utilized to assess the language alignment within LLMs as well as the comprehension and generation capabilities across different languages. Multilingual interaction involves multitask language understanding using multiple languages, which can be employed to evaluate the multilingual knowledge and reasoning abilities of LLMs.
4.1.1 Multilingual Translation
We employ multilingual translation to assess the multilingual alignment within LLMs, which entails producing outputs that retain the same meaning but in different languages. We conduct evaluation on the Flores-101 and WMT22 benchmarks. For metrics, BLEU (sacrebleu) (Post, 2018) and COMET (Rei et al., 2022) are used to assess the quality of LLMs’ translation. BLEU score measures the statistical similarity based on n-gram accuracy, COMET score measures the semantic similarity using cross-lingual pre-trained models, which is currently regarded as the most human-aligned evaluation metric for translation tasks.
Table 1: Mulitlingual translation preformance on WMT22 benchmark. X indicates other 100 languages in Flores-101, and the results are averaged over these 100 languages.
[…]
BLEU COMET BLEU COMET BLEU COMET BLEU COMET
Llama-1-7B 14.07 60.94 6.93 49.73 0.93 40.88 1.85 44.88
BayLing-1-7B 14.70 61.93 7.04 49.33 1.58 46.22 1.56 48.78
Llama-2-7B-Chat 15.39 63.95 7.45 50.97 1.75 47.19 1.57 45.70
BayLing-2-7B 17.71 67.15 8.02 52.37 2.70 51.44 2.32 49.37
Llama-3-8B-Instruct 25.20 76.60 16.59 67.17 11.79 71.91 8.95 63.57
BayLing-2-8B 26.77 77.03 17.91 70.88 11.31 69.43 10.64 67.86
[…]
Flores-101 Flores-101 benchmark encompasses 101 languages from around the world, and the sentences is sourced from various domains, including news, travel guides and books. Due to the rarity of some low-resource languages, LLMs may suffer from off-target issues. To address this, we adopt a 1-shot setting (i.e., randomly selecting an example from the dev set) to help LLMs follow the target language through in-context learning. We compare BayLing models with their corresponding foundational models, and the results are shown in Table 1.
The results in Table 1 indicate that BayLing achieves better performance in most translation directions between 100 languages and Chinese/English. Specifically, compared to the foundation LLMs Llama-1-7B and Llama-2-7B-Chat, which have relatively weak multilingual capabilities, BayLing effectively scales their language understanding and generation capabilities to over 100 languages, leading to significantly improved translation performance. Furthermore, Figures 6(a), 6(b), 7(a) and 7(b) illustrate the specific BLEU score improvements achieved by BayLing across 100 languages. BayLing consistently delivers the highest translation quality for most languages, particularly in translation directions to low-resource languages. This demonstrates BayLing’s potential to enhance LLM in serving such low-source linguistic communities.
Table 2: Mulitlingual translation preformance on WMT22 benchmark. The bold and underlined results indicate the first and second best, respectively.
Systems $\RightarrowZh En\RightarrowDe$ $\RightarrowCs En\RightarrowJa$ $\RightarrowRu En\RightarrowUk$
COMET BELU COMET BELU COMET BELU COMET BELU COMET BELU COMET BELU
closed-sourced
GPT-4 87.49 43.98 87.44 35.38 90.77 34.53 89.87 24.71 88.87 30.45 88.46 26.71
GPT-3.5-turbo 86.81 44.99 86.93 34.12 90.05 32.71 83.26 22.22 87.52 29.59 87.43 25.87
Google Translate 87.34 49.89 87.08 38.27 91.28 48.10 88.64 26.50 88.91 35.04 88.63 32.05
open-sourced
Llama-2-7B-Chat 67.90 17.50 72.22 16.74 65.17 11.69 69.66 9.52 67.60 12.47 66.94 10.65
Llama-2-13B-Chat 75.23 24.31 77.25 20.35 75.42 16.18 78.46 13.56 77.19 17.11 75.41 14.75
Vicuna-7B-v1.5 81.40 29.54 75.25 16.65 71.84 13.63 74.80 11.28 77.66 17.95 74.96 13.26
Vicuna-13B-v1.5 84.01 34.69 81.99 24.22 77.97 17.47 85.45 17.54 83.31 21.60 81.32 17.86
Llama-3-8B-Instruct 80.55 30.10 82.18 25.83 83.24 23.41 65.43 10.57 82.92 23.53 80.69 18.88
BayLing-1-7B 84.43 38.19 82.18 25.66 76.85 15.64 71.23 4.51 74.72 14.85 76.01 11.66
BayLing-1-13B 84.62 37.92 82.69 25.62 78.22 16.43 71.39 6.05 71.01 12.77 66.83 8.32
BayLing-2-7B 85.94 39.71 82.76 25.65 80.54 17.81 84.94 16.43 82.03 19.72 75.85 12.25
BayLing-2-13B 86.65 42.87 83.79 26.61 82.93 18.52 83.60 15.79 85.23 22.95 80.89 14.60
BayLing-2-8B 85.75 41.49 84.53 29.59 87.55 24.57 83.34 16.82 86.79 26.41 85.97 21.81
[…]
COMET BELU COMET BELU COMET BELU COMET BELU COMET BELU COMET BELU
closed-sourced
GPT-4 82.79 27.20 85.62 33.87 87.43 48.67 83.20 24.57 86.18 43.51 85.67 40.47
GPT-3.5-turbo 82.64 26.13 85.47 32.94 86.75 45.99 82.39 22.14 85.95 41.79 85.32 39.00
Google Translate 80.81 28.63 84.75 33.21 86.95 49.26 81.69 23.17 84.81 43.54 85.55 41.60
open-sourced
Llama-2-7B-Chat 75.31 15.42 80.14 24.53 78.18 28.91 74.11 11.38 80.56 31.23 79.41 28.68
Llama-2-13B-Chat 75.90 16.45 81.43 25.69 81.28 33.97 76.37 13.37 81.60 32.72 81.19 31.55
Vicuna-7B-v1.5 75.42 16.80 79.07 23.57 76.34 24.46 72.13 10.89 78.63 27.95 78.29 25.74
Vicuna-13B-v1.5 78.47 19.41 83.25 29.19 81.71 34.51 75.22 13.66 82.18 33.74 82.54 33.03
Llama-3-8B-Instruct 80.44 21.57 83.84 29.37 83.47 39.49 78.83 17.00 84.08 37.05 82.75 34.53
BayLing-1-7B 77.48 20.31 83.19 28.16 82.03 35.98 72.16 11.63 82.48 34.74 81.38 33.07
BayLing-1-13B 77.72 20.12 83.02 27.34 81.65 33.87 72.14 12.23 82.07 33.95 81.41 32.67
BayLing-2-7B 79.07 22.09 82.56 27.33 80.66 31.91 76.35 15.12 81.19 29.37 80.34 28.60
BayLing-2-13B 79.47 23.43 83.07 28.81 81.63 34.08 76.55 14.95 81.71 31.75 80.56 30.09
BayLing-2-8B 79.75 22.58 83.55 28.99 83.16 37.43 79.25 18.61 83.15 34.07 81.89 31.98
[…]
Improving Mulitlingual Generation Capabilities We have observed that foundational models often exhibit off-target issues when generating low-resource languages. In contrast, BayLing demonstrates significantly enhanced multilingual generation capabilities, consistently improving translation performance from English to other languages. This indicates that BayLing can activate the multilingual generation abilities of LLMs solely through cross-lingual translation data, without the need for extensive multilingual instruction data. This finding is crucial for efficiently enhancing the multilingual capabilities of LLMs, as it is nearly impossible to collect instruction data covering more than 100 languages while multilingual translation data is relatively abundant and easier to obtain. BayLing’s approach of transferring generation capabilities from high-resource to low-resource languages through language alignment offers an efficient solution for enhancing the multilingual generation capabilities of LLMs.
The superior multilingual translation capabilities on Flores-101 and WMT22 underscores BayLing’s potential as a leading tool in the field of multilingual translation, offering significant advancements in multilingual capabilities of LLM.
4.1.2 Multilingual Multi-task Evaluation
We assessed the multilingual performance of BayLing using several benchmarks. All evaluations were conducted through the Language Model Evaluation Harness [^9] https://github.com/EleutherAI/lm-evaluation-harness (Gao et al., 2023), an open-source, unified framework designed to assess LLMs across a wide variety of evaluation tasks. Each result was obtained in a zero-shot setting. The models Llama-2-7B, Llama-2-7B-Chat, Llama-3-8B-Instruct, Vicuna-7B and Mistral-7B served as baselines for comparison. The multilingual benchmarks are discribed as follows.
Belebele (Bandarkar et al., 2023) Belebele is a multiple-choice machine reading comprehension benchmark, which evaluates mono- and multi-lingual models across different resource levels with rigorously checked questions. Each question has four multiple-choice answers and is linked to a short passage from the FLORES-200 dataset.
Multilingual HellaSwag (Dac Lai et al., 2023) Multilingual HellaSwag is a multilingual adaptation of HellaSwag, a benchmark dataset designed to assess commonsense inference. Despite its questions being straightforward for humans, state-of-the-art models struggle with it, highlighting the challenges in AI comprehension.
XNLI (Conneau et al., 2018) XNLI is an evaluation dataset created by extending the MultiNLI corpus to multiple languages, including low-resource ones like Swahili and Urdu. It serves as a standardized benchmark for assessing cross-lingual sentence understanding, aiming to foster research in this area.
Multilingual ARC (Dac Lai et al., 2023) The Multilingual ARC, a multilingual extension of ARC (Clark et al., 2018), encompasses science examination queries, stratified into a Challenge Set comprising intricate questions and an Easy Set. All queries adhere to a multiple-choice structure.
(a) Belebele
(b) Multilingual HellaSwag
(c) XNLI
(d) Multilingual ARC
Figure 8: Multilingual multi-task performance of BayLing on low-resource languages.
Figure 8(a), 8(b), 8(c), 8(d) provide detailed illustrations of the experimental outcomes on the Belebele, Multilingual HellaSwag, XNLI, Multilingual ARC benchmarks across several low-resource languages. The BayLing-2-7B and BayLing-2-8B models demonstrate notable performance benefits. Among these, BayLing-2-8B consistently delivers the best results across most of the low-resource languages evaluated. Meanwhile, BayLing-2-7B outperforms other 7B models in most of these languages. Remarkably, BayLing-2-7B even surpasses the Llama-3-8B-Instruct model in the Swati language subset on Belebele benchmark (belebele_ssw_Latn).
Note that BayLing’s training data does not include instruction data for these low-resource languages but only cross-lingual instructions between these low-resource languages and Chinese/English. Therefore, the performance improvements observed in these low-resource languages demonstrate that BayLing effectively transfers knowledge and understanding capability from high-resource languages to low-resource ones through language alignment. Overall, BayLing’s approach of leveraging language alignment for capability transfer offers an efficient solution to enhance LLM’s performance in low-resource languages.
4.2 General Capability
Furthermore, we evaluated the general capability of BayLing on the following benchmarks employing the same settings as described in section 4.1.2.
CMMLU (Li et al., 2023) CMMLU serves as a specialized evaluation benchmark tailored to assess the knowledge and reasoning capacities of LLMs in within the context of Chinese language and culture. Encompassing a wide range of subjects, CMMLU includes 67 topics, which vary from basic to advanced professional levels.
C-Eval (Huang et al., 2023) C-Eval is an exhaustive Chinese evaluation suite designed for foundation models. It features a total of 13,948 multiple-choice questions, covering 52 distinct disciplines across four levels of difficulty.
Arabic EXAMS (Hardalov et al., 2020) The Arabic EXAMS comprise the Arabic segment of EXAMS, a resource dedicated to multilingual high school examination questions. This section includes five subjects: Islamic Studies, Biology, Physics, General Science, and Social Studies.
ANLI (Nie et al., 2020) Adversarial NLI (ANLI) is a dataset assembled through an iterative adversarial procedure involving both human and model participation. It is structured into three rounds, each escalating in difficulty and complexity. Additionally, each question-answer pair in the dataset is supplemented with explanations provided by the annotators.
CB (De Marneffe et al., 2019; Wang et al., 2019) CB (CommitmentBank) is a corpus featuring texts with embedded clauses evaluated for the author’s commitment to their truth. This corpus is used in a three-class textual entailment task. Examples are organized with a premise and a corresponding hypothesis extracted from the embedded clause.
GLUE (Wang et al., 2018) The GLUE benchmark is a benchmark for evaluating natural language understanding systems. It consists of nine language understanding tasks and a diagnostic dataset for assessing model performance across linguistic phenomena.
ACLUE (Zhang and Li, 2023) ACLUE is a benchmark designed to evaluate large language models’ comprehension of ancient Chinese, featuring 15 tasks across multiple domains. The questions, covering historical periods from the Xia to the Ming dynasty, are presented in a multiple-choice format.
GSM8K (Cobbe et al., 2021) GSM8K is a benchmark used to evaluate the math capability of LLMs as described in section 4.1.2.
Figure 9: Scores on general benchmarks.
Figure 9 illustrates the performance of BayLing on the specified benchmarks. BayLing-2-7B and BayLing-2-8B demonstrate exceptional performance across several benchmarks. Notably, BayLing-2-8B outperforms all other models, achieving a score of 0.8214 on the CommitmentBank Benchmark. BayLing-2-7B attains the highest performance on GLUE and GSM8K benchmarks. Despite not being specifically trained for it, BayLing-2-7B and BayLing-2-8B still deliver comparable performances on other benchmarks when compared to other models. Overall, BayLing enhances the multilingual capabilities of LLMs, especially in low-resource languages, without significantly impacting the performance in high-resource languages. This indicates that BayLing effectively mitigates multilingual conflicts within LLM through language alignment.
Figure 10: Effect of language alignment on multilingual benchmark Belebele.
Figure 11: Effect of language alignment on Chinese and English general tasks.
4.3 Effect of Language Alignment
To validate the effect of language alignment brought by cross-lingual instructions, we conducted an ablation study on cross-lingual instructions. Specifically, we removed all cross-lingual instructions from the training data, denoting this variant as BayLing-2-8B (w/o cross-lingual instructions).
Improving Performance of Low-resource Languages Figure 10 compares the performance of BayLing-2-8B and BayLing-2-8B (w/o cross-lingual instructions) on the multilingual benchmark Belebele. The results show that cross-lingual instructions significantly enhance LLM performance in low-resource languages. This indicates that cross-lingual instructions successfully help LLM achieve language alignment, thereby transferring knowledge and comprehension from high-resource languages to low-resource ones. When removing all cross-lingual instructions, the performance of LLMs in low-resource languages is adversely affected due to catastrophic forgetting. Therefore, involving cross-lingual instructions in supervised fine-tuning is both efficient and crucial for improving the multilingual capabilities of LLMs.
Avoiding Inter-language Conflicts Figure 11 compares BayLing-2-8B and BayLing-2-8B (w/o cross-lingual instructions) on the Chinese/English benchmark. When removing cross-lingual instructions, we observed a significant performance decline in Chinese benchmark, indicating LLM will suffer from conflicts between Chinese and English instructions. The presence of numerous cross-lingual instructions between Chinese and English largely prevents these conflicts. Therefore, to simultaneously enhance LLM performance across multiple languages, introducing cross-lingual instructions is an effective way to avoid inter-language conflicts.
5 Conclusion
In this study, we develop BayLing 2, which enhances LLM’s multilingual capabilities through language alignment. Adhering to an efficiency-focused approach, BayLing 2 transfers knowledge and generative abilities from high-resource languages to low-resource languages within LLM via language alignment. Comprehensive evaluation results demonstrate that BayLing 2 achieves outstanding translation performance across over 100 languages, possesses superior multilingual knowledge and understanding capability, and maintains robust proficiency in high-resource languages of Chinese and English.
Acknowledgements
We extend our heartfelt gratitude to everyone who contributed to the development of BayLing 2. In particular, we would like to thank Ms. Xiaohong Wang for her insightful feedback and valuable suggestions regarding the use of OneAiNexus, as well as for her exceptional support in organizing resources, providing computational infrastructure, and facilitating the presentation of BayLing 2. Besides, we would like to express thanks to the Sothis.AI for the support in training of BayLing 2. Special thanks are due to the team of Nanjing Institute of InforSuperBahn - Intelligent Computility Platform Research Center, who played an indispensable role in maintaining the computational resources, designing the BayLing 2’s webpage and demonstrating the system interface.
Appendix A Flores-101 Benchmark
Table 3, 4, 5, and 6 report the numerical results of Llama-3-8B-Instruct and BayLing-2-8B on the Flores-101 benchmark.
Table 3: BLEU scores of Llama-3-8B-Instruct on Flores-101 benchmark.
Llama-3-8B-Instruct
[…]
afr Afrikaans 50.65 37.33 16.48 14.22 lug Ganda 6.08 1.11 3.87 0.44
amh Amharic 2.63 0.44 1.33 0.35 luo Luo 4.01 0.89 2.95 0.35
ara Arabic 33.28 19.63 14.38 11.31 mal Malayalam 20.44 4.72 9.56 2.51
asm Assamese 16.17 7.20 7.99 4.60 mar Marathi 22.72 10.64 10.65 6.35
ast Asturian 38.63 29.44 15.56 14.74 mkd Macedonian 38.47 27.85 16.29 13.82
azj North Azerbaijani 19.25 10.84 11.36 7.77 mlt Maltese 38.54 24.08 13.22 10.92
bel Belarusian 19.74 14.90 12.78 10.03 mon Mongolian 12.46 3.93 7.67 2.48
ben Bengali 24.80 12.00 12.65 7.34 mri Maori 12.61 7.70 5.80 4.20
bos Bosnian 39.23 25.86 17.27 12.11 msa Malay 39.26 32.11 16.46 14.87
bul Bulgarian 37.27 30.83 16.51 15.50 mya Burmese 7.42 1.21 3.76 0.86
cat Catalan 43.40 39.56 17.80 19.91 nld Dutch 31.91 28.18 16.72 16.28
ceb Cebuano 27.00 16.21 11.27 7.21 nob Norwegian Bokmål 41.49 29.28 16.35 13.48
ces Czech 38.47 28.96 17.05 15.01 npi Nepali 23.40 11.74 11.28 6.62
ckb Central Kurdish 16.17 4.80 7.02 2.72 nso Pedi 6.63 1.96 4.14 0.68
cym Welsh 39.84 24.66 12.15 10.29 nya Nyanja 7.10 1.87 4.36 0.70
dan Danish 46.38 37.87 17.31 15.82 oci Occitan 45.82 27.01 16.16 11.99
deu German 41.65 34.96 18.64 17.45 orm Oromo 2.07 0.79 1.96 0.47
ell Modern Greek 32.57 22.41 15.73 13.05 ory Odia 15.71 1.90 7.48 0.99
est Estonian 29.61 18.09 13.96 9.55 pan Panjabi 22.23 7.08 10.35 3.70
fas Persian 31.33 23.02 14.44 13.59 pol Polish 29.08 22.77 15.96 13.99
fin Finnish 31.78 21.84 15.44 12.16 por Portuguese 47.99 45.68 18.29 20.59
fra French 43.17 45.32 18.48 23.99 pus Pushto 18.50 4.94 8.78 2.68
ful Fulah 3.58 0.90 3.04 0.31 ron Romanian 39.91 34.53 17.38 16.94
gle Irish 25.45 13.90 9.81 6.66 rus Russian 33.95 28.81 17.56 15.56
glg Galician 39.13 32.84 17.26 17.32 slk Slovak 35.54 23.77 16.33 11.63
guj Gujarati 21.76 6.02 9.93 3.30 slv Slovenian 31.65 21.68 15.10 11.33
hau Hausa 13.07 8.13 5.86 3.45 sna Shona 6.07 1.63 4.30 0.76
heb Hebrew 36.86 22.01 15.48 11.05 snd Sindhi 17.58 8.96 7.34 4.89
hin Hindi 31.88 24.48 14.61 14.04 som Somali 6.36 2.77 4.05 1.21
hrv Croatian 35.45 24.16 16.36 11.33 spa Spanish 31.50 28.89 17.05 18.16
hun Hungarian 31.88 24.10 16.31 13.72 srp Serbian 39.61 30.24 17.08 13.97
hye Armenian 29.80 6.04 13.60 3.69 swe Swedish 46.13 39.44 18.00 16.02
ibo Igbo 11.61 5.07 5.99 2.25 swh Swahili 29.48 16.48 11.17 6.90
ind Indonesian 39.34 37.60 17.87 18.27 tam Tamil 18.64 5.49 9.65 3.29
isl Icelandic 24.72 12.43 10.56 6.52 tel Telugu 24.19 6.29 10.36 3.59
ita Italian 33.37 30.87 18.14 18.37 tgk Tajik 20.40 12.22 10.35 7.74
jav Javanese 23.77 10.94 10.33 4.27 tgl Tagalog 37.24 24.10 14.55 12.48
jpn Japanese 23.08 24.79 14.00 16.27 tha Thai 25.58 18.41 14.39 12.31
kam Kamba 4.55 1.54 3.55 0.58 tur Turkish 31.84 23.84 16.02 13.54
kan Kannada 20.75 4.91 10.45 2.98 ukr Ukrainian 37.24 29.18 16.25 14.37
kat Georgian 20.37 3.90 11.11 2.45 umb Umbundu 3.28 0.55 2.34 0.35
kaz Kazakh 22.91 11.61 12.06 7.63 urd Urdu 24.72 13.71 11.57 8.00
kea Kabuverdianu 26.56 6.66 9.84 1.68 uzb Uzbek 20.70 18.13 10.65 12.28
khm Khmer 12.44 1.05 6.09 0.84 vie Vietnamese 32.72 33.91 16.43 21.28
kir Kirghiz 15.58 10.19 8.63 7.01 wol Wolof 4.83 1.30 3.75 0.57
kor Korean 25.41 18.78 15.47 12.69 xho Xhosa 8.60 1.76 5.08 0.67
lao Lao 6.58 0.29 3.35 0.10 yor Yoruba 6.53 2.71 3.90 1.68
lav Latvian 27.95 17.82 13.37 9.50 zho_simpl Chinese 26.07 22.53 - -
lin Lingala 5.85 2.54 4.28 1.08 zho_trad traditional Chinese 24.48 19.46 21.63 23.55
lit Lithuanian 27.90 16.88 13.47 9.96 zul Zulu 8.11 1.87 4.42 0.80
ltz Luxembourgish 33.72 18.60 13.03 8.78 eng English - - 22.53 26.07
Table 4: BLEU scores of BayLing-2-8B on Flores-101 benchmark.
BayLing-2-8B
[…]
afr Afrikaans 51.55 38.76 15.83 15.83 lug Ganda 9.41 3.59 4.49 1.88
amh Amharic 4.97 1.87 1.67 1.31 luo Luo 8.00 5.15 4.33 2.80
ara Arabic 35.36 21.19 13.54 12.06 mal Malayalam 16.22 3.29 6.15 2.15
asm Assamese 16.12 7.79 6.72 5.75 mar Marathi 26.48 12.81 12.10 7.92
ast Asturian 39.22 29.93 15.66 15.26 mkd Macedonian 38.95 27.18 15.46 14.92
azj North Azerbaijani 21.91 9.39 9.06 7.51 mlt Maltese 40.28 27.29 13.10 13.89
bel Belarusian 23.08 13.24 10.39 9.17 mon Mongolian 14.69 6.68 7.99 4.61
ben Bengali 23.14 11.66 9.94 7.59 mri Maori 15.75 17.43 7.02 12.28
bos Bosnian 38.90 24.94 16.50 13.83 msa Malay 40.46 34.66 17.02 17.27
bul Bulgarian 37.79 27.42 15.69 15.68 mya Burmese 5.10 1.30 2.28 0.94
cat Catalan 44.45 37.33 18.06 20.48 nld Dutch 33.71 26.77 16.46 17.00
ceb Cebuano 30.29 21.80 11.87 12.49 nob Norwegian Bokmål 41.70 26.69 16.06 13.61
ces Czech 38.26 25.71 16.48 14.33 npi Nepali 28.46 17.17 12.59 10.55
ckb Central Kurdish 17.03 7.53 7.23 4.81 nso Pedi 12.21 9.03 5.51 5.08
cym Welsh 39.66 27.38 13.08 12.24 nya Nyanja 12.68 6.29 6.65 4.07
dan Danish 45.98 35.50 16.81 17.36 oci Occitan 47.85 33.78 16.54 14.79
deu German 43.25 34.12 18.26 17.26 orm Oromo 5.31 2.21 2.21 1.71
ell Modern Greek 34.58 20.24 14.15 13.18 ory Odia 8.35 1.67 2.70 1.06
est Estonian 30.49 16.11 10.91 9.48 pan Panjabi 19.06 7.57 6.79 4.65
fas Persian 34.21 24.49 15.00 15.42 pol Polish 30.97 20.72 14.87 12.97
fin Finnish 31.50 19.24 13.30 11.89 por Portuguese 47.89 45.74 18.93 22.73
fra French 44.31 45.36 18.63 24.60 pus Pushto 22.64 6.56 9.20 4.44
ful Fulah 7.25 2.46 4.14 1.57 ron Romanian 42.63 33.28 17.37 18.93
gle Irish 27.75 16.25 10.19 9.14 rus Russian 35.61 27.68 15.20 16.66
glg Galician 41.44 34.02 17.67 18.25 slk Slovak 35.78 21.30 14.28 12.15
guj Gujarati 18.35 6.67 7.74 4.42 slv Slovenian 32.12 19.69 13.59 11.87
hau Hausa 19.81 10.06 7.24 6.10 sna Shona 11.68 5.34 5.59 3.41
heb Hebrew 34.79 21.79 13.56 11.73 snd Sindhi 23.93 17.00 8.69 10.17
hin Hindi 33.85 25.45 14.79 15.84 som Somali 12.52 5.38 5.02 3.05
hrv Croatian 35.55 22.53 15.33 13.07 spa Spanish 34.75 29.33 17.04 18.26
hun Hungarian 32.34 21.10 13.28 14.18 srp Serbian 39.53 28.33 15.58 15.32
hye Armenian 18.19 5.07 4.92 3.58 swe Swedish 45.93 36.58 18.09 17.93
ibo Igbo 16.35 10.73 7.26 7.28 swh Swahili 29.56 19.19 11.36 9.83
ind Indonesian 40.64 38.12 17.50 20.80 tam Tamil 16.94 5.73 7.49 3.56
isl Icelandic 26.87 12.68 9.95 6.81 tel Telugu 18.23 5.96 7.40 3.82
ita Italian 35.91 29.22 17.88 17.76 tgk Tajik 21.77 15.30 9.56 11.52
jav Javanese 29.19 20.48 10.22 11.73 tgl Tagalog 39.94 28.51 14.44 15.80
jpn Japanese 27.97 25.64 16.37 18.73 tha Thai 28.93 17.47 14.08 11.62
kam Kamba 8.16 3.50 4.39 2.11 tur Turkish 33.43 22.28 14.63 13.78
kan Kannada 14.49 4.75 6.13 3.40 ukr Ukrainian 37.41 26.95 15.65 15.59
kat Georgian 11.06 2.88 4.36 1.86 umb Umbundu 5.82 4.22 2.95 2.87
kaz Kazakh 23.63 13.89 10.49 9.51 urd Urdu 26.40 13.53 12.30 8.70
kea Kabuverdianu 34.60 29.40 12.31 14.48 uzb Uzbek 24.39 17.99 11.43 12.49
khm Khmer 14.07 2.19 6.55 1.86 vie Vietnamese 35.23 34.89 15.61 22.46
kir Kirghiz 18.47 12.66 8.01 9.56 wol Wolof 9.28 4.34 4.45 2.87
kor Korean 27.60 18.40 15.43 13.80 xho Xhosa 13.14 4.08 5.79 2.43
lao Lao 10.52 1.82 5.15 1.10 yor Yoruba 12.04 9.19 4.98 7.74
lav Latvian 29.48 16.30 12.03 9.82 zho_simpl Chinese 30.24 25.61 - -
lin Lingala 11.61 11.95 5.93 7.70 zho_trad traditional Chinese 28.00 20.25 17.82 19.60
lit Lithuanian 28.03 16.12 11.98 9.67 zul Zulu 12.84 6.59 5.70 4.26
ltz Luxembourgish 35.92 22.59 12.47 12.57 eng English - - 25.61 30.24
Table 5: COMET scores of Llama-3-8B-Instruct on Flores-101 benchmark.
Llama-3-8B-Instruct
[…]
afr Afrikaans 85.79 82.75 79.07 76.71 lug Ganda 49.86 39.62 48.30 40.12
amh Amharic 57.47 38.74 50.99 35.71 luo Luo 46.09 34.60 45.78 36.63
ara Arabic 84.85 77.90 79.64 75.42 mal Malayalam 82.12 54.48 74.40 49.07
asm Assamese 78.26 62.26 72.23 56.98 mar Marathi 82.20 56.38 75.18 49.97
ast Asturian 80.99 70.31 78.66 68.53 mkd Macedonian 85.40 82.46 79.96 77.19
azj North Azerbaijani 82.72 74.79 76.83 68.66 mlt Maltese 75.13 62.43 68.83 59.95
bel Belarusian 80.32 77.39 78.85 74.53 mon Mongolian 73.22 53.00 65.69 49.14
ben Bengali 84.66 67.34 78.66 61.77 mri Maori 57.72 56.14 54.45 54.72
bos Bosnian 86.51 85.48 81.80 81.65 msa Malay 86.33 86.71 80.29 82.04
bul Bulgarian 85.60 84.17 80.74 79.12 mya Burmese 73.03 50.27 65.28 46.07
cat Catalan 86.74 85.53 82.64 81.52 nld Dutch 85.72 85.80 82.13 82.25
ceb Cebuano 72.62 62.99 67.83 60.00 nob Norwegian Bokmål 86.63 86.87 81.59 82.31
ces Czech 86.80 86.06 81.91 83.05 npi Nepali 84.40 65.83 77.16 56.32
ckb Central Kurdish 67.45 69.63 62.18 67.04 nso Pedi 49.40 42.12 48.67 43.00
cym Welsh 81.50 71.46 72.45 64.23 nya Nyanja 51.77 39.38 50.41 39.27
dan Danish 88.59 86.51 82.44 81.05 oci Occitan 81.76 66.45 77.59 63.47
deu German 87.66 85.67 83.25 81.05 orm Oromo 48.00 44.67 46.05 47.08
ell Modern Greek 85.45 82.65 81.00 77.82 ory Odia 79.47 47.13 71.51 41.06
est Estonian 84.71 77.30 78.32 72.13 pan Panjabi 82.41 56.27 75.00 50.09
fas Persian 85.51 80.16 80.68 76.40 pol Polish 84.37 85.39 81.77 82.69
fin Finnish 87.88 85.98 81.46 80.16 por Portuguese 87.90 88.06 83.34 83.48
fra French 87.57 86.28 83.00 81.79 pus Pushto 76.71 50.35 71.00 43.86
ful Fulah 46.30 36.09 46.03 37.68 ron Romanian 86.94 86.82 81.63 80.42
gle Irish 75.46 60.43 68.96 56.50 rus Russian 84.90 85.59 82.30 81.43
glg Galician 86.28 84.41 82.33 81.94 slk Slovak 85.45 80.06 80.76 77.14
guj Gujarati 83.51 59.21 74.24 52.37 slv Slovenian 84.16 79.80 79.63 77.13
hau Hausa 63.46 59.62 57.57 55.83 sna Shona 50.90 38.50 50.16 39.94
heb Hebrew 85.42 75.54 79.71 71.34 snd Sindhi 76.17 49.99 67.98 42.34
hin Hindi 86.14 72.12 80.65 63.87 som Somali 54.83 44.04 50.93 43.12
hrv Croatian 85.54 84.45 81.08 81.00 spa Spanish 85.64 85.28 83.57 82.58
hun Hungarian 85.98 84.73 81.21 77.73 srp Serbian 85.98 83.67 80.73 78.84
hye Armenian 85.44 58.56 78.90 55.00 swe Swedish 88.08 88.20 82.94 80.34
ibo Igbo 58.27 53.70 55.43 50.65 swh Swahili 78.61 69.99 70.94 63.60
ind Indonesian 87.30 89.21 81.77 85.19 tam Tamil 80.96 60.47 74.26 55.75
isl Icelandic 78.18 63.75 72.37 60.57 tel Telugu 82.73 56.46 74.54 51.16
ita Italian 86.16 86.76 83.61 83.25 tgk Tajik 69.92 68.34 65.52 66.27
jav Javanese 75.31 70.76 68.50 61.76 tgl Tagalog 83.36 77.46 76.79 73.13
jpn Japanese 86.03 88.26 82.65 85.59 tha Thai 85.91 83.39 82.02 79.59
kam Kamba 48.40 33.91 47.96 34.32 tur Turkish 86.43 83.47 80.18 78.05
kan Kannada 82.02 54.19 75.12 49.32 ukr Ukrainian 85.49 85.64 81.01 80.66
kat Georgian 81.79 49.97 76.46 47.32 umb Umbundu 45.68 33.53 44.83 37.23
kaz Kazakh 82.37 70.70 76.29 64.64 urd Urdu 83.29 67.81 77.05 62.28
kea Kabuverdianu 70.07 56.13 66.16 46.47 uzb Uzbek 80.19 78.27 72.90 72.81
khm Khmer 74.15 45.33 67.68 41.98 vie Vietnamese 85.35 85.97 82.55 83.92
kir Kirghiz 78.34 64.65 71.78 59.11 wol Wolof 48.08 37.41 47.81 40.19
kor Korean 85.64 84.84 81.85 80.24 xho Xhosa 55.13 39.41 52.34 38.15
lao Lao 60.88 32.31 53.76 30.66 yor Yoruba 50.57 46.30 48.48 46.88
lav Latvian 82.45 71.96 77.13 68.26 zho_simpl Chinese 85.40 85.12 - -
lin Lingala 50.05 40.91 50.11 41.12 zho_trad traditional Chinese 84.82 85.70 87.43 90.30
lit Lithuanian 82.24 74.78 78.12 72.44 zul Zulu 55.38 42.01 52.00 40.05
ltz Luxembourgish 73.87 52.23 70.23 50.93 eng English - - 85.12 85.40
Table 6: COMET scores of BayLing-2-8B on Flores-101 benchmark.
BayLing-2-8B
[…]
afr Afrikaans 87.38 82.77 76.93 76.33 lug Ganda 52.34 54.41 47.80 55.35
amh Amharic 58.46 49.90 49.78 46.53 luo Luo 49.37 53.25 47.44 53.08
ara Arabic 85.49 81.47 76.61 77.17 mal Malayalam 76.81 53.88 64.05 49.89
asm Assamese 75.40 61.88 66.47 57.93 mar Marathi 84.00 61.50 74.95 55.19
ast Asturian 82.33 68.74 76.10 65.80 mkd Macedonian 85.77 83.44 77.39 77.94
azj North Azerbaijani 83.50 71.42 71.40 68.93 mlt Maltese 76.80 63.82 68.15 60.42
bel Belarusian 81.56 75.77 72.15 72.44 mon Mongolian 74.76 63.14 66.60 59.34
ben Bengali 82.05 67.65 72.81 62.06 mri Maori 60.55 65.82 56.43 65.98
bos Bosnian 86.69 85.46 79.17 82.11 msa Malay 87.13 87.12 79.21 81.73
bul Bulgarian 86.31 84.48 77.62 80.05 mya Burmese 60.32 48.38 53.58 45.40
cat Catalan 87.42 85.18 80.95 81.70 nld Dutch 86.50 85.82 79.37 82.29
ceb Cebuano 74.56 66.23 66.35 63.70 nob Norwegian Bokmål 87.63 86.18 78.41 81.41
ces Czech 87.01 86.20 79.42 81.39 npi Nepali 86.33 72.79 77.50 64.69
ckb Central Kurdish 68.04 72.53 60.30 70.81 nso Pedi 53.76 55.76 50.90 55.37
cym Welsh 82.80 73.36 72.70 65.56 nya Nyanja 56.96 54.35 54.57 54.77
dan Danish 88.60 87.16 80.00 82.53 oci Occitan 82.98 66.67 76.59 62.64
deu German 88.43 85.42 80.85 79.83 orm Oromo 51.98 61.08 46.52 60.07
ell Modern Greek 86.30 83.03 77.55 78.65 ory Odia 67.58 47.06 54.55 43.06
est Estonian 85.04 77.37 72.89 72.54 pan Panjabi 78.43 57.99 65.19 52.88
fas Persian 86.17 82.74 79.21 78.08 pol Polish 84.65 85.58 77.17 81.26
fin Finnish 87.62 85.71 76.73 80.74 por Portuguese 88.37 88.53 82.17 84.36
fra French 88.14 86.43 81.70 81.82 pus Pushto 78.81 57.44 69.19 52.47
ful Fulah 50.38 49.69 47.84 49.27 ron Romanian 88.17 87.47 79.59 82.31
gle Irish 77.62 64.82 68.62 61.00 rus Russian 85.49 87.13 77.67 83.61
glg Galician 87.43 85.06 81.31 80.94 slk Slovak 85.69 80.35 76.51 76.29
guj Gujarati 79.51 61.78 68.02 55.96 slv Slovenian 84.95 79.18 76.07 75.20
hau Hausa 70.44 68.11 61.42 65.61 sna Shona 56.12 53.40 52.40 53.55
heb Hebrew 84.52 77.60 75.31 73.13 snd Sindhi 78.95 60.37 68.54 52.39
hin Hindi 87.48 74.74 78.84 66.93 som Somali 62.89 59.67 53.83 57.87
hrv Croatian 86.11 85.12 78.05 81.40 spa Spanish 86.58 85.55 81.64 82.25
hun Hungarian 86.22 84.28 75.17 80.15 srp Serbian 85.94 83.72 78.32 80.11
hye Armenian 76.46 54.09 60.19 51.67 swe Swedish 88.92 88.34 80.68 83.90
ibo Igbo 62.72 65.76 58.49 65.44 swh Swahili 79.48 72.72 71.24 68.74
ind Indonesian 87.77 89.74 80.27 85.85 tam Tamil 76.36 61.98 66.96 58.39
isl Icelandic 80.19 65.65 68.99 60.63 tel Telugu 77.62 56.91 66.75 51.67
ita Italian 86.98 86.88 81.82 82.94 tgk Tajik 70.08 71.18 62.96 71.06
jav Javanese 78.81 81.17 68.05 76.92 tgl Tagalog 84.71 79.28 74.49 74.54
jpn Japanese 86.79 87.94 83.12 85.10 tha Thai 86.19 82.18 79.48 77.51
kam Kamba 51.18 48.26 49.49 48.68 tur Turkish 87.04 83.07 77.55 77.89
kan Kannada 74.36 53.53 63.26 49.64 ukr Ukrainian 85.79 86.42 77.66 83.00
kat Georgian 68.95 44.91 58.06 43.45 umb Umbundu 48.02 52.10 45.96 52.57
kaz Kazakh 82.51 74.61 71.55 70.50 urd Urdu 83.03 69.03 75.18 63.64
kea Kabuverdianu 75.72 64.54 68.23 61.44 uzb Uzbek 82.38 78.14 72.98 74.18
khm Khmer 71.97 50.53 65.81 48.50 vie Vietnamese 86.15 86.53 79.57 83.87
kir Kirghiz 80.57 72.88 69.29 68.85 wol Wolof 51.85 56.18 48.48 57.08
kor Korean 86.17 84.66 80.73 80.95 xho Xhosa 59.23 55.73 53.95 55.86
lao Lao 63.71 44.71 57.43 42.45 yor Yoruba 57.90 64.74 51.41 65.55
lav Latvian 83.60 72.25 72.71 68.30 zho_simpl Chinese 61.08 53.89 - -
lin Lingala 56.47 60.44 53.36 59.81 zho_trad traditional Chinese 85.56 86.38 80.02 88.93
lit Lithuanian 82.79 74.51 73.92 71.44 zul Zulu 59.43 62.17 53.89 61.08
ltz Luxembourgish 76.32 51.80 68.17 51.12 eng English - - 85.92 86.22
Appendix B Language Code of Multilingual Benchmarks
Table 7 reports the language codes for the low-resource languages in Figure 8.
Table 7: Language code of Figure 8.
Language Code Language
belebele_bam_Latn Bambara
belebele_ben_Latn Bengali
belebele_hau_Latn Hausa
belebele_ilo_Latn Ilocano
belebele_kin_Latn Kinyarwanda
belebele_lao_Laoo Lao
belebele_lin_Latn Lingala
belebele_lug_Latn Luganda
belebele_luo_Latn Luo
belebele_nso_Latn Northern Sotho
belebele_nya_Latn Chichewa (Nyanja)
belebele_pbt_Arab Pashto
belebele_plt_Latn Plateau Malagasy
belebele_sna_Latn Shona
belebele_sot_Latn Southern Sotho
belebele_ssw_Latn Swazi
belebele_tso_Latn Tsonga
belebele_yor_Latn Yoruba
belebele_zul_Latn Zulu
hellaswag_eu Basque (Euskara)
hellaswag_mr Marathi
hellaswag_ne Nepali
hellaswag_vi Vietnamese
xnli_bg Bulgarian
xnli_de German
xnli_es Spanish
xnli_sw Swahili
xnli_ur Urdu
xnli_zh Chinese
arc_ar Arabic
arc_eu Basque (Euskara)
arc_hi Hindi
arc_hy Armenian
arc_mr Marathi
arc_ne Nepali
arc_sr Serbian
arc_uk Ukrainian
arc_zh Chinese
Appendix C Numerical Results of General Capability
Table 8 reports the numerical results of general capability in Figure 9. Table 9, 10, 11, and 12 report the numerical results of multilingual benchmarks in Figure 8.
Table 8: Numerical results of general capability benchmarks.
Model cmmlu ceval-valid aexams ammlu anli
(knowledge) (knowledge) (knowledge) (knowledge) (comprehension)
Llama-2-7B 0.2726 0.2972 0.2272 0.2717 0.3697
Llama-2-7B-Chat 0.3257 0.3239 0.2272 0.2624 0.4106
Llama-3-8B-Instruct 0.5120 0.5111 0.3520 0.3838 0.4634
Vicuna-7B-v1.5 0.3514 0.3603 0.2439 0.2962 0.3847
Mistral-7B 0.3825 0.3990 0.2495 0.2763 0.3800
BayLing-1-7B 0.3136 0.3046 0.2346 0.2458 0.3622
BayLing-2-7B 0.3916 0.3908 0.2607 0.2900 0.4419
BayLing-2-8B 0.4839 0.4643 0.3371 0.3440 0.4372
Model cb glue aclue gsm8k
(comprehension) (comprehension) (comprehension) (math)
Llama-2-7B 0.4464 0.4271 0.2747 0.0000
Llama-2-7B-Chat 0.6071 0.4863 0.2755 0.0000
Llama-3-8B-Instruct 0.8036 0.5877 0.3963 0.0265
Vicuna-7B-v1.5 0.7143 0.4729 0.2976 0.0000
Mistral-7B 0.5000 0.5144 0.3192 0.0000
BayLing-1-7B 0.5000 0.5411 0.2763 0.0000
BayLing-2-7B 0.7321 0.6158 0.3228 0.1531
BayLing-2-8B 0.8214 0.5559 0.3272 0.0311
Table 9: Numerical results of Belebele benchmark.
Model belebele_bam_Latn belebele_ben_Latn belebele_hau_Latn belebele_ilo_Latn belebele_kin_Latn
Llama-2-7B 0.2278 0.2422 0.2322 0.2244 0.2167
Llama-2-7B-Chat 0.2278 0.2244 0.2333 0.2422 0.2289
Llama-3-8B-Instruct 0.3033 0.3189 0.3478 0.3533 0.3167
Vicuna-7B-v1.5 0.2411 0.2533 0.2478 0.2667 0.2600
Mistral-7B 0.2756 0.2878 0.3022 0.2878 0.3144
BayLing-1-7B 0.2578 0.2856 0.2367 0.2300 0.2778
BayLing-2-7B 0.2889 0.2800 0.3000 0.3133 0.2700
BayLing-2-8B 0.3189 0.3500 0.3533 0.3589 0.3356
Model belebele_lin_Latn belebele_lug_Latn belebele_luo_Latn belebele_nso_Latn belebele_nya_Latn
Llama-2-7B 0.2233 0.2356 0.2244 0.2378 0.2267
Llama-2-7B-Chat 0.2167 0.2456 0.2433 0.2322 0.2078
Llama-3-8B-Instruct 0.3178 0.3144 0.3044 0.3067 0.2833
Vicuna-7B-v1.5 0.2400 0.2533 0.2433 0.2500 0.2522
Mistral-7B 0.2789 0.3122 0.2967 0.2689 0.2856
BayLing-1-7B 0.2611 0.2744 0.2767 0.2556 0.2733
BayLing-2-7B 0.2822 0.2678 0.2822 0.2978 0.2711
BayLing-2-8B 0.3411 0.3433 0.3122 0.3122 0.3167
Model belebele_plt_Latn belebele_sna_Latn belebele_sot_Latn belebele_ssw_Latn belebele_tso_Latn
Llama-2-7B 0.2444 0.2156 0.2378 0.2311 0.2189
Llama-2-7B-Chat 0.2278 0.2244 0.2533 0.2433 0.2500
Llama-3-8B-Instruct 0.3422 0.3244 0.3133 0.2889 0.3211
Vicuna-7B-v1.5 0.2589 0.2556 0.2711 0.2633 0.2533
Mistral-7B 0.3044 0.2944 0.3044 0.2856 0.2833
BayLing-1-7B 0.2644 0.2622 0.2389 0.2522 0.2433
BayLing-2-7B 0.2744 0.2867 0.2633 0.300 0.2900
BayLing-2-8B 0.3711 0.3333 0.3400 0.3244 0.3267
Model belebele_zul_Latn belebele_lao_Laoo belebele_pbt_Arab belebele_yor_Latn
Llama-2-7B 0.2256 0.2178 0.2144 0.2200
Llama-2-7B-Chat 0.2489 0.2344 0.2233 0.2300
Llama-3-8B-Instruct 0.3244 0.3011 0.3411 0.2944
Vicuna-7B-v1.5 0.2711 0.2567 0.2689 0.2533
Mistral-7B 0.2956 0.2967 0.2900 0.2667
BayLing-1-7B 0.2633 0.2267 0.2344 0.2578
BayLing-2-7B 0.2967 0.2856 0.3022 0.2456
BayLing-2-8B 0.3378 0.3656 0.3078 0.3233
Table 10: Numerical results of multilingual HellaSwag benchmark.
Model hellaswag_eu hellaswag_mr hellaswag_ne hellaswag_vi
Llama-2-7B 0.2588 0.2618 0.2634 0.3522
Llama-2-7B-Chat 0.2594 0.2599 0.2627 0.3445
Llama-3-8B-Instruct 0.2759 0.2770 0.2791 0.4021
Vicuna-7B-v1.5 0.2595 0.2644 0.2659 0.3556
Mistral-7B 0.2633 0.2617 0.2733 0.3614
BayLing-1-7B 0.2577 0.2576 0.2598 0.2831
BayLing-2-7B 0.2585 0.2631 0.2645 0.3458
BayLing-2-8B 0.2765 0.2807 0.2948 0.4025
Table 11: Numerical results of XNLI benchmark.
Model xnli_bg xnli_de xnli_es xnli_sw xnli_ur xnli_zh
Llama-2-7B 0.4249 0.4699 0.4064 0.3478 0.3357 0.3639
Llama-2-7B-Chat 0.3723 0.4321 0.3956 0.3410 0.3410 0.3699
Llama-3-8B-Instruct 0.4518 0.4932 0.4863 0.3622 0.3454 0.4052
Vicuna-7B-v1.5 0.4177 0.4663 0.4763 0.3446 0.3398 0.4056
Mistral-7B 0.4309 0.4639 0.4406 0.3510 0.3414 0.3960
BayLing-1-7B 0.4213 0.4474 0.4558 0.3450 0.3438 0.3614
BayLing-2-7B 0.3594 0.4036 0.3936 0.3345 0.3422 0.3731
BayLing-2-8B 0.4510 0.4972 0.4908 0.3655 0.3454 0.4096
Table 12: Numerical results of multilingual ARC benchmark.
Model arc_ar arc_eu arc_hi arc_hy arc_mr arc_ne arc_sr arc_uk arc_zh
Llama-2-7B 0.0128 0.0079 0.0283 0.0036 0.0208 0.0214 0.0470 0.0565 0.0556
Llama-2-7B-Chat 0.0034 0.0035 0.0128 0.0018 0.0078 0.0060 0.0180 0.0359 0.0214
Llama-3-8B-Instruct 0.1172 0.0360 0.0839 0.0245 0.0545 0.0496 0.1138 0.1437 0.1248
Vicuna-7B-v1.5 0.0197 0.0053 0.0317 0.0036 0.0130 0.0111 0.0393 0.0582 0.0855
Mistral-7B 0.0188 0.0097 0.0068 0.0009 0.0035 0.0043 0.0753 0.0804 0.0632
BayLing-1-7B 0.0145 0.0062 0.0214 0.0018 0.0173 0.0137 0.0359 0.0325 0.0752
BayLing-2-7B 0.0205 0.0053 0.0146 0.0027 0.0069 0.0120 0.0248 0.0214 0.0718
BayLing-2-8B 0.1223 0.0518 0.0865 0.0255 0.0580 0.0667 0.1172 0.1600 0.1684